Calculate the pH of a solution that is 0.650 M HOCl and 1.10 M OCl-.
Answer #1

Ka of HOCl = 3.5*10^-8

pKa = - log (Ka)

= - log(3.5*10^-8)

= 7.456

use:

pH = pKa + log {[conjugate base]/[acid]}

= 7.456+ log {1.1/0.65}

= 7.684

Answer: 7.68